# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.1.3 Existence
Taking the above subsequence $\big{\{}(\varphi^{k},\mu^{k},\sigma^{k})\big{\}}$ that is convergent on $[0,1]$, using now the uniform estimates on the time interval $[0,2]$, we can exact a further convergent subsequence with the limit $(\varphi_{(2)},\mu_{(2)},\sigma_{(2)})$, which yields a weak solution to problem (3.1)–(3.3) on $[0,2]$. From the construction, it holds $(\varphi_{(2)},\mu_{(2)},\sigma_{(2)})=(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})$ on $[0,1]$. By successive extractions, for any $N\in\mathbb{Z}^{+}$, we are able to obtain a weak solution $(\varphi_{(N)},\mu_{(N)},\sigma_{(N)})$ to problem (3.1)–(3.3) on $[0,N]$ such that $(\varphi_{(N+1)},\mu_{(N+1)},\sigma_{(N+1)})=(\varphi_{(N)},\mu_{(N)},\sigma_{(N)})$ on $[0,N]$. Hence, applying the diagonal extraction process as in [8], we can find a subsequence $\big{\{}(\varphi^{k},\mu^{k},\sigma^{k})\big{\}}$ that converges towards a weak solution of problem (3.1)–(3.3) on any interval $[0,T]$, the limit of this subsequence is indeed a global weak solution defined on $[0,+\infty)$, as in the statement of Theorem 3.1-(1). ∎
3.1.4 Conditional uniqueness
Different from the single convective Cahn–Hilliard equation studied in [1, 2], uniqueness of weak solutions to problem (3.1)–(3.3) in three dimensions is unknown because of the low regularity of $\bm{v}$ and $\sigma$. On the other hand, for the approximate solution $(\varphi^{k},\mu^{k},\sigma^{k})$, we observed that $\sigma^{k}$ can be used as a test function in its weak formulation such that
@@ 3.1.4 Conditional uniqueness
For the convergent subsequence $\big{\{}(\varphi^{k},\mu^{k},\sigma^{k})\big{\}}$ constructed in the previous section, we can follow the standard argument in [8, Chapter V, Section 1.4] and pass to the limit as $k\to+\infty$ to conclude the following energy inequality for the weak solution:
$\frac{1}{2}\|\sigma(t)\|^{2}+\int_{0}^{t}\|\nabla\sigma(s)\|^{2}\,\mathrm{d}s\leq\frac{1}{2}\|\sigma_{0}\|^{2}+\chi\int_{0}^{t}\!\int_{\Omega}\nabla\sigma\cdot\nabla\varphi\,\mathrm{d}x\mathrm{d}s,\quad\forall\,t>0.$ (3.31)
@@
Thanks to the additional regularity $\sigma_{2}\in L^{8}(0,T;L^{4}(\Omega))$, for any $\xi\in L^{2}(0,T;H^{1}(\Omega))$, we have
@@
which implies that $\partial_{t}\sigma_{2}\in L^{2}(0,T;(H^{1}(\Omega))^{\prime})$ by comparison in the weak formulation for $\sigma_{2}$. In a similar manner, for any $\xi\in L^{8}(0,T;L^{4}(\Omega))$, it holds
$\left|\int_{0}^{T}\!\int_{\Omega}(\bm{v}\cdot\nabla\sigma_{1})\xi\,\mathrm{d}x\mathrm{d}s\right|$
$\quad\leq\int_{0}^{T}\|\bm{v}\|_{\bm{L}^{4}}\|\nabla\sigma_{1}\|\|\xi\|_{L^{4}}\mathrm{d}s\leq\int_{0}^{T}\|\bm{v}\|^{\frac{1}{4}}\|\nabla\bm{v}\|^{\frac{3}{4}}\|\nabla\sigma_{1}\|\|\xi\|_{L^{4}}\,\mathrm{d}s$
$\quad\leq C\|\bm{v}\|_{L^{\infty}(0,T;\bm{L}^{2}(\Omega))}^{\frac{1}{4}}\|\bm{v}\|_{L^{2}(0,T;\bm{H}^{1}(\Omega))}^{\frac{3}{4}}\|\sigma_{1}\|_{L^{2}(0,T;H^{1}(\Omega))}\|\xi\|_{L^{8}(0,T;L^{4}(\Omega))},$
which further implies that $\partial_{t}\sigma_{1}\in\big{(}L^{8}(0,T;L^{4}(\Omega))\big{)}^{\prime}$.
The above observations allow us to test the weak formulations for $\sigma_{1}$, $\sigma_{2}$ by $\sigma_{2}$, respectively. Meanwhile, $\sigma_{1}$ can be used as a test function in the weak formulation of $\sigma_{2}$ as well. Integrating the resultants on $[0,t]\subset[0,T]$, we obtain
@@
$(\sigma_{1}(t),\sigma_{2}(t))+\int_{0}^{t}\!\int_{\Omega}\bm{v}\cdot\nabla(\sigma_{1}\sigma_{2})\,\mathrm{d}x\mathrm{d}s+2\int_{0}^{t}\!\int_{\Omega}\nabla\sigma_{1}\cdot\nabla\sigma_{2}\,\mathrm{d}x\mathrm{d}s$
@@
On the other hand, we recall that the weak solution $(\varphi_{1},\sigma_{1})$ satisfies the energy inequality (3.31) on $[0,t]\subset[0,T]$. Adding it with (3.32) and subtracting (3.33), we find the following estimate on the difference $\sigma$:
@@
The term on the right-hand side can be estimated by Poincaré’s inequality and Young’s inequality such that
@@
Thanks to (H2), we can handle $(\mu,\varphi)$ by a similar argument like that in [27, Section 4]:
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\varphi\|_{V_{0}^{-1}}^{2}+\frac{1}{2}\|\nabla\varphi\|^{2}\leq\frac{1}{2}\|\sigma\|^{2}+C\big{(}1+\|\bm{v}\|_{\bm{H}^{1}}^{2}\big{)}\|\varphi\|_{V_{0}^{-1}}^{2}.$
Integrating on $[0,t]\subset[0,T]$, we obtain
@@
If $\chi=0$, we infer from (3.34) that $\sigma(t)=0$ for $t\in[0,T]$. Then by (3.35) and Gronwall’s lemma, we can conclude $\varphi(t)=0$ for $t\in[0,T]$. Otherwise, if $\chi\neq 0$, from (3.34) and (3.35), we find
@@
which together with Gronwall’s lemma again yields $\varphi(t)=\sigma(t)=0$ for $t\in[0,T]$.

### v2

@@ 3.1.3 Existence
Taking the above subsequence $\big\{(\varphi^{k},\mu^{k},\sigma^{k})\big\}$ that is convergent on $[0,1]$, using now the uniform estimates on the time interval $[0,2]$, we can exact a further convergent subsequence with the limit $(\varphi_{(2)},\mu_{(2)},\sigma_{(2)})$, which yields a weak solution to problem (3.1)–(3.3) on $[0,2]$. From the construction, we find that
$(\varphi_{(2)},\mu_{(2)},\sigma_{(2)})=(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})\quad\text{on}\ \ [0,1].$
By successive extractions, for any $N\in\mathbb{Z}^{+}$, we are able to obtain a weak solution $(\varphi_{(N)},\mu_{(N)},\sigma_{(N)})$ to problem (3.1)–(3.3) on $[0,N]$ such that
$(\varphi_{(N+1)},\mu_{(N+1)},\sigma_{(N+1)})=(\varphi_{(N)},\mu_{(N)},\sigma_{(N)})\quad\text{on}\ \ [0,N].$
Hence, applying the diagonal extraction process as in [10], we can find a subsequence $\big\{(\varphi^{k},\mu^{k},\sigma^{k})\big\}$ that converges towards a weak solution of problem (3.1)–(3.3) on any interval $[0,T]$, the limit of this subsequence is indeed a global weak solution defined on $0,+\infty)$.
Let $\big\{(\varphi^{k},\mu^{k},\sigma^{k})\big\}$ be the convergent subsequence constructed above. We observed that $\sigma^{k}$ can be used as a test function in its weak formulation such that
@@
Applying the standard argument in [10, Chapter V, Section 1.4] and passing to the limit as $k\to+\infty$, we can conclude the energy inequality (3.5). This completes the proof of Theorem 3.1-(1). ∎
3.1.4 Conditional uniqueness
Unlike the single convective Cahn–Hilliard equation studied in [1, 3], the uniqueness of weak solutions to problem (3.1)–(3.3) in three dimensions is not known due to the low regularity of $\bm{v}$ and $\sigma$. However, with the help of the energy inequality (3.5), we are able to establish a result about conditional uniqueness.
@@ 3.1.4 Conditional uniqueness
Thanks to the additional regularity $\sigma_{2}\in L^{8}(0,T;L^{4}(\Omega))$ and the interpolation inequality $\|\bm{v}\|_{\bm{L}^{4}}\leq C\|\bm{v}\|^{\frac{1}{4}}\|\nabla\bm{v}\|^{\frac{3}{4}}$, for any $\xi\in L^{2}(0,T;H^{1}(\Omega))$, we have
@@
which implies that $\partial_{t}\sigma_{2}\in L^{2}(0,T;(H^{1}(\Omega))^{\prime})$ by comparison in the weak formulation for $\sigma_{2}$ (note that the regularity of $\varphi_{2}$ is sufficient for a test function in $L^{2}(0,T;H^{1}(\Omega))$). In a similar manner, for any $\zeta\in L^{8}(0,T;L^{4}(\Omega))$, it holds
$\left|\int_{0}^{T}\!\!\int_{\Omega}(\bm{v}\cdot\nabla\sigma_{1})\zeta\,\mathrm{d}x\mathrm{d}s\right|$
$\quad\leq\int_{0}^{T}\|\bm{v}\|_{\bm{L}^{4}}\|\nabla\sigma_{1}\|\|\zeta\|_{L^{4}}\mathrm{d}s\leq\int_{0}^{T}\|\bm{v}\|^{\frac{1}{4}}\|\nabla\bm{v}\|^{\frac{3}{4}}\|\nabla\sigma_{1}\|\|\zeta\|_{L^{4}}\,\mathrm{d}s$
$\quad\leq C\|\bm{v}\|_{L^{\infty}(0,T;\bm{L}^{2}(\Omega))}^{\frac{1}{4}}\|\bm{v}\|_{L^{2}(0,T;\bm{H}^{1}(\Omega))}^{\frac{3}{4}}\|\sigma_{1}\|_{L^{2}(0,T;H^{1}(\Omega))}\|\zeta\|_{L^{8}(0,T;L^{4}(\Omega))},$
which implies that
$\partial_{t}\sigma_{1}\in(L^{8}(0,T;L^{4}(\Omega)))^{\prime}+L^{2}(0,T;(H^{1}(\Omega))^{\prime}).$
The above observations allow us to test the weak formulations for $\sigma_{1}$, $\sigma_{2}$ by $\sigma_{2}$, respectively. Meanwhile, $\sigma_{1}$ can be used as a test function in the weak formulation of $\sigma_{2}$ as well. Integrating the resultants on $0,t]\subset[0,T]$, we obtain
@@
$(\sigma_{1}(t),\sigma_{2}(t))+\,\underbrace{\int_{0}^{t}\!\int_{\Omega}\bm{v}\cdot\nabla(\sigma_{1}\sigma_{2})\,\mathrm{d}x\mathrm{d}s}_{=0}\,+2\int_{0}^{t}\!\int_{\Omega}\nabla\sigma_{1}\cdot\nabla\sigma_{2}\,\mathrm{d}x\mathrm{d}s$
@@
On the other hand, we recall that the weak solution $(\varphi_{1},\sigma_{1})$ satisfies the energy inequality (3.5) on $0,t]\subset[0,T]$. Adding it with (3.29) and subtracting (3.30), we arrive at the following estimate on the difference $\sigma$:
@@
The term on the right-hand side can be estimated by Poincaré’s inequality, the Poincaré–Wirtinger inequality and Young’s inequality such that
@@
Thanks to (H2), we can handle $(\mu,\varphi)$ by a similar argument like that in [31, Section 4]:
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\varphi\|_{V_{0}^{-1}}^{2}+\frac{1}{2}\|\nabla\varphi\|^{2}\leq\frac{1}{2}\|\sigma\|^{2}+C\big(1+\|\bm{v}\|_{\bm{H}^{1}}^{2}\big)\|\varphi\|_{V_{0}^{-1}}^{2}.$ (3.32)
Integrating on $0,t]\subset[0,T]$, we obtain
@@
If $\chi=0$, we infer from (3.31) and Gronwall’s lemma that $\sigma(t)=0$ for $t\in[0,T]$. Then by (3.33) and Gronwall’s lemma, we can further conclude $\varphi(t)=0$ for $t\in[0,T]$. Otherwise, if $\chi\neq 0$, from (3.31) and (3.33), we find
@@
which together with Gronwall’s lemma again yields that $\varphi(t)=\sigma(t)=0$ for $t\in[0,T]$.